OpenStudy (anonymous):

i bought two lovebirds and later sold them, at a loss for 24$ each. if my percent loss on each lovebird was numerically equal to the dollar price i paid for that love bird, what were the two dollar amounts i originally paid for the love birds?

OpenStudy (anonymous):

$48?

OpenStudy (triciaal):

@H20Delirious don't just guess answers help the "asker" to get to the correct solution

OpenStudy (triciaal):

cost - %loss = sold sold for $24 each so sold for $48 my percent loss on each lovebird was numerically equal to the dollar price i paid for that love bird, cost = c and % means per 100 c -(loss/c)*100 =24 loss/c%= c loss = c^2/10 100 - loss = 24 cost 76 sorry